OpenStudy (mony01):

Verify by differentiation and substitution that the given function is a solution. y' = y^2 + 4 (l t l < pi/4); y = 2 tan 2t

OpenStudy (anonymous):

Sorry I don't know how to do this stuff ;s You can go in the math chat and ask for help.

Miracrown (miracrown):

The directions say to differentiate and substitute You want to differentiate the given y = 2tan(2t) solution Any ideas on how to take the derivative of that?

Miracrown (miracrown):

It's function inside of a function so what derivative rule do we need? @mony01

OpenStudy (mony01):

is it the chain rule?

OpenStudy (mony01):

@Miracrown

myininaya (myininaya):

What is the derivative of the inside, 2t?

Miracrown (miracrown):

The easiest way to handle this problem is to first find the derivative, y', assuming the given y is a solution. Would you like to work the derivative, y', while we observe? @mony01

OpenStudy (mony01):

would the derivative be 2(2tan2t)*2

OpenStudy (mony01):

@Miracrown

OpenStudy (alekos):

first find the derivative of tant then multiply this by the derivative of 2t (chain rule), then multiply the lot by 2

OpenStudy (alekos):

that will give you the LHS

OpenStudy (mony01):

@alekos the derivative of tan t is sec^2 t right?

OpenStudy (dumbcow):

yes

OpenStudy (mony01):

I know that since y=2tan2t I need to plug this in into the other equation right so it would be y'= (2tan2t)^2 +4 so now is it y'= (2sec^2(2t))^2 +4

OpenStudy (dumbcow):

\[y' = 2*(2\sec^2 (2t)) = 4 \sec^2 (2t)\] \[y' = y^2 + 4\] \[\rightarrow 4 \sec^2(2t) = (2 \tan(2t))^2 + 4\]

OpenStudy (mony01):

where did you put the +4 when you got 4sec^2(2t)?

OpenStudy (dumbcow):

you are confusing 2 different things the differential equation ---> y' = y^2 + 4 and y' as simply the derivative of y, which is given as 2tan(2t) they are different but equal, that is why we can substitute it into the diff equation

OpenStudy (mony01):

oh so the whole point is to solve for y' using y only

OpenStudy (dumbcow):

yes ... they gave you what the solution "y" is to verify you have to plug it into the diff equation but you also have to plug in something for y' thus taking derivative of y gives y' no solving at all really

OpenStudy (mony01):

oh thank you i understand now so we didn't even have to deal with the +4 right

OpenStudy (dumbcow):

well its still part of the diff equation but it has nothing to do with y'

OpenStudy (mony01):

ok thank you so much

OpenStudy (dumbcow):

yw :)